OpenStudy (awkwardpanda):

1. what is an equation of the parabola with vertex at the origin and focus (-6 0)?

OpenStudy (anonymous):

you know what this looks like?

OpenStudy (anonymous):

i ask because you need that to start

OpenStudy (awkwardpanda):

hmmm. a graph??

OpenStudy (anonymous):

yeah, the graph

OpenStudy (anonymous):



OpenStudy (anonymous):

looks like that so you know a) it opens to the left b) the \(y\) term is squared

OpenStudy (anonymous):

since the vertex is \((0,0)\) it is going to look like \[y^2=-4px\] where \(p\) is the distance between the vertex and the origin that distance is pretty clearly \(6\)

OpenStudy (anonymous):

let me know when you get \[y^2=-24x\]

OpenStudy (anonymous):

we can check it if you like

OpenStudy (anonymous):

yup looks good to me http://www.wolframalpha.com/input/?i=parabola+y^2%3D-24x

OpenStudy (awkwardpanda):

so it would be \[y= \frac{ 1 }{ 24} x^{2}\]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

guess i was not clear the \(y\) term is squared, not the \(x\) term

OpenStudy (anonymous):

also there is a minus sign

OpenStudy (awkwardpanda):

WAIT SO ITS -24 Y^2

OpenStudy (anonymous):

lol no

OpenStudy (anonymous):

\[y^2=-24x\] really trust me

OpenStudy (awkwardpanda):

deng it. ugh i meant -1/24 y^2 i have answer choices

OpenStudy (anonymous):

yeah, multiple choice the death of clear thinking and mathematics no matter, if you see \[-\frac{1}{24}y^2=x\] go with that one

OpenStudy (awkwardpanda):

would the 1 be negative.

OpenStudy (anonymous):

i don't know dear a minus sign is a minus sign it does not live in the top or the bottom of a fraction right?

OpenStudy (awkwardpanda):

because none of my answer choices have that. it's either positive 1 over 24 or negative 1 over 24

OpenStudy (anonymous):

if you want to post your answer choices i will pick one out for you

OpenStudy (anonymous):

go with \(\frac{-1}{24}\)

OpenStudy (awkwardpanda):

yeah that one seems right ^ thank you for being patient :D

OpenStudy (anonymous):

which is i hope pretty clearly identical to \(-\frac{1}{24}\)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

you made me blush, i have to be patient